eviant:

PLS HELP!!! You drop a ball from a height of 1.5 meters. Each curved path has 75% of the height of the previous path. a. Write a rule for the sequence using centimeters. The initial height is given by the term n = 1. b. What height will the ball be at the top of the sixth path?

Vocaloid:

a. since it wants everything in cm, first convert 1.5m to cm. this will be your initial value. then it's (initial value)*(0.75)^n

eviant:

(150)(0.75)^n

eviant:

@Vocaloid

Vocaloid:

good then for b) simply plug in n = 6 to find the height on the sixth bounce

eviant:

675

Vocaloid:

hm. you may have made a mistake somewhere. since the ball height is getting smaller we would expect the height to be much less than 150cm not greater (150)(0.75)^6 = ?

Vocaloid:

don't forget the exponent ^

eviant:

26.6?

Vocaloid:

almost, it's 26.696 so it would round up to 26.7 cm

eviant:

A(n) = 150 ∙ (75)x − 1; 355,957,031,250 cm A(n) = 1.5 ∙ (0.75)x − 1; 0.36 cm A(n) = 150 ∙ (0.75)x − 1; 35.6 cm A(n) = 0.75 ∙ (1.5)x − 1; 5.7 cm

Vocaloid:

oh duh I'm silly since the index starts at n = 1 we have to change the formula to account for this by subtracting 1 from the index so the formula is 150(0.75)^(n-1) (they use x for some reason)

Vocaloid:

so plugging in n = 6 into this new formula gives like 36ish cm as the result which is closer to choice C